# A Set of Tutorials for the LAMMPS Simulation Package [Article v1.0]

### v1

@@ 3.8 Tutorial 8: Reactive Molecular Dynamics
The goal of this tutorial is to create a model of a carbon nanotube (CNT) embedded in a polymer melt made of polystyrene (PS) (Fig. 47). The REACTER protocol is used to simulate the polymerization of styrene monomers, and the polymerization reaction is followed in time [67, 15, 68]. In contrast with AIREBO (Tutorial 2) and ReaxFF (Tutorial 5), the REACTER protocol relies on the use of a classical force field.
[…]
@@ 3.8.1 Creating the system
The class2 styles compute a 6/9 Lennard-Jones potential [69]. The class2 bond, angle, dihedral, and improper styles are used as well, see the documentation for a description of their respective potentials. The mix sixthpower imposes the following mixing rule for the calculation of the cross coefficients:
@@
[…]
@@
[…]
@@
The fix halt command is used to stop the box shrinkage once the target density is reached.
@@
[…]
@@
For visualization purposes, the atoms from the CNT group is moved to the center of the box using fix recenter. As the time progresses, the system density, $\rho$, gradually converges toward the target value of $0.8$ g/cm3 (Fig. 48 a). Meanwhile, the total energy of the system initially evolves rapidly, reflecting the densification process, and then eventually stabilizes (Fig. 48 b). The final state is shown in Fig. 47.
@@ 3.8.3 Simulating the reaction
Here, the read__\__data command is used to import the previously generated mixing.data file. All other commands have been introduced in earlier parts of the tutorial.
@@
With the stabilization keyword, the bond/react command will stabilize the atoms involved in the reaction using the nve/limit command with a maximum displacement of $0.03\,\mathrm{\AA{}}$. By default, each reaction is stabilized for 60 time steps. Each react keyword corresponds to a reaction, e.g., a transformation of mol1 into mol2 based on the atom map M-M.rxnmap. Implementation details about each reaction, such as the reaction distance cutoffs and the frequency with which to search for reaction sties, are also specified in this command.
@@
Add the following commands to polymerize.lmp to operate in the NVT ensemble while ensuring that the CNT remains centered in the simulation box:
@@
Here, the thermo custom command is used to print the cumulative reaction counts from fix rxn. Run the simulation using LAMMPS. As the simulation progresses, polymer chains are observed forming (Fig. 49). During this reaction process, the temperature of the system remains well-controlled (Fig. 50 a), while the number of reactions, $N_{r}$, increases with time (Fig. 50 b).
@@
S.G. conceived and wrote the original online tutorials and underlying Sphinx documentation for lammpstutorials.github.io. J.G. is the principal author of fix bond/react and type labels support in LAMMPS. He revised the tutorials to incorporate type labels and wrote Tutorial 8. A.K. developed the LAMMPS–GUI software and assisted in revising the tutorials for use with it. All authors participated in the revision and finalization of the manuscript.
@@
S.G. acknowledges funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement $\text{N}^{\circ}\;101065060$. A.K. acknowledges financial support by Sandia National Laboratories under POs 2149742 and 2407526.
@@ A.1 Installation
Precompiled versions of LAMMPS–GUI are available for Linux, macOS, and Windows on the LAMMPS GitHub Release page [21]. The Linux version is provided in two formats: as compressed tar archive (.tar.gz) and as a Flatpak bundle [70]. The macOS version is distributed as a .dmg installer image, while the Windows version comes as an executable installer package.
@@
Download the archive (e.g., LAMMPS-Linux-x86_64-GUI-29Aug2024_update2.tar.gz) and unpack it. This will create a folder named LAMMPS–GUI containing the included commands, which can be launched directly using “./lammps-gui” or “./lmp”, for example. Adding this folder to the PATH environment variable will make these commands accessible from everywhere, without the need for the “./” prefix.
@@ A.1.2 Installing the Linux Flatpak Bundle
You have to have Flatpak support installed on Linux machine to be able to use the Flatpak bundle. Download the bundle file (e.g., LAMMPS-Linux-x86_64-GUI-29Aug2024_update2.flatpak) and then install it using the following command:
@@
LAMMPS-Linux-x86_64-GUI-29Aug2024_update2.flatpak
@@ A.1.3 Installing the macOS Application Bundle
After downloading the macOS app bundle image file (e.g., LAMMPS-macOS-multiarch-GUI-29Aug2024_update2.dmg), double-click on it. In the dialog that opens drag the LAMMPS–GUI app bundle into the Applications folder. To enable command-line access, follow the instructions in the README.txt file. These macOS app-bundles contain native executables for both, Intel and Apple CPUs.
@@ A.1.4 Installing the Windows package
Download the LAMMPS–GUI installer for Windows (e.g., LAMMPS-Win10-64bit-GUI-29Aug2024_update2.exe). Windows may warn you that the file is from an unknown developer and was downloaded from the internet. This happens because neither the installer nor the LAMMPS–GUI application (or any other included applications) have been cryptographically signed. You will need to choose to keep the file, and when launching the installer, confirm that you want to run it despite the warning.
@@ A.2 Opening, Editing, and Saving Files
Files can also be opened from the “File” menu. You can select a file through a dialog and then open it. Additionally, a history of the last five opened files is maintained, with entries to open them directly. Finally, the Ctrl-O keyboard shortcut can also be used to open a file.
When integrated into a desktop environment, it is also possible to open files with a “.lmp” extension or use drag-and-drop.
@@
If the file in the editor has unsaved changes, the word “modified” will appear in the window title. The current input buffer can be saved by selecting “Save” or “Save As…” from the “File” menu. You can also click the “Save” icon on the left side of the status bar, or use the Ctrl-S keyboard shortcut.
@@ Appendix B Running LAMMPS on the Command-Line without the GUI
See Ref. 71 for a complete description on how to run LAMMPS.

### v2

@@ 3.8 Tutorial 8: Reactive Molecular Dynamics
The goal of this tutorial is to create a model of a carbon nanotube (CNT) embedded in a polymer melt made of polystyrene (PS) (Fig. 47). The REACTER protocol is used to simulate the polymerization of styrene monomers, and the polymerization reaction is followed in time [68, 15, 69]. In contrast with AIREBO (Tutorial 2) and ReaxFF (Tutorial 5), the REACTER protocol relies on the use of a classical force field that does not inherently model bond formation or breaking, but instead couples with an external algorithm to simulate polymerization reactions.
Figure 47: Initial configuration for Tutorial 8. The system consists of 200 styrene molecules packed around a single-walled CNT, with a mass density for the whole system of $0.9\leavevmode\nobreak\ \text{g/cm}^{3}$.
@@ 3.8.1 Creating the system
The class2 styles compute a 6/9 Lennard-Jones potential [70]. The class2 bond, angle, dihedral, and improper styles are used as well, see the documentation for a description of the respective potential form they, each, prescribe. The tail yes option adds long-range van der Waals tail corrections to the energy and pressure. The mix sixthpower imposes the following mixing rule for the calculation of the cross coefficients:
@@
The CNT is approximately $1.1\leavevmode\nobreak\ \text{nm}$ in diameter and $1.6\,\text{nm}$ in length, oriented along the $x$-axis. The simulation box is initially 12.0 nm in the two other dimensions before densification, making it straightforward to fill the box with styrene. To add 200 styrene molecules to the simulation box, we will use the styrene.mol molecule template file. Include the following commands to mixing.lmp:
@@
These commands should be familiar from previous tutorials.
Then, let us densify the system to a target value of $0.9\leavevmode\nobreak\ \text{g/cm}^{3}$ by imposing the shrinking of the simulation box at a constant rate. The dimension parallel to the CNT axis is maintained fixed because the CNT is periodic in that direction. Add the following commands to mixing.lmp:
@@
The fix halt command is used to stop the box shrinkage once the target density is reached, and the other commands should be familiar from previous tutorials.
@@
For the following $10\leavevmode\nobreak\ \text{ps}$, let us equilibrate the densified system in the constant-volume ensemble, and write the final state of the system in a file named mixing.data:
@@
For visualization purposes, the atoms of the CNT group are moved to the center of the box using fix recenter. As the time progresses, the system density, $\rho$, gradually converges toward the target value of $0.9$ g/cm3 (Fig. 48 a). Meanwhile, the total energy of the system initially evolves rapidly, reflecting the densification process, and then eventually stabilizes (Fig. 48 b). The final state is shown in Fig. 47.
@@ 3.8.2 Reaction templates
Note 3.50.
The data stored in molecule templates include atom coordinates, partial charges, molecule IDs, atom types, and interaction types for bonds, angles, dihedrals and impropers. The map files contain information about the reaction. The first mandatory section of the map files begins with the keyword “InitiatorIDs” and lists the two atom IDs of the initiator atom pair in the pre-reacted molecule template. The second mandatory section begins with the keyword “Equivalences” and lists a one-to-one correspondence between atom IDs of the pre- and post-reacted templates. Some atoms in the pre-reacted template that are not reacting may have missing topology with respect to the simulation. For example, the pre-reacted template may contain an atom that, in the simulation, is currently connected to the rest of a long polymer chain. These are referred to as edge atoms, and are also specified in the map file in the “EdgeIDs” section.
@@ 3.8.3 Simulating the reaction
Here, the read_data command is used to import the previously generated mixing.data file. All other commands have been introduced in earlier parts of the tutorial.
@@
With the stabilization keyword, the fix bond/react command will stabilize the atoms involved in the reaction using the fix nve/limit command with a maximum displacement of $0.03\,\mathrm{\AA {}}$. The fix nve/limit command functions similar to fix nve, but restricts how far atoms can move in a single time step, even with very large forces. By default, each reaction is stabilized for 60 time steps. Each react keyword corresponds to a reaction, e.g., a transformation of mol1 into mol2. Implementation details about each reaction, such as the reaction distance cutoffs and the frequency with which to search for reaction sites, are also specified in this command.
@@
Add the following commands to polymerize.lmp to carry out the dynamics using a Nosé-Hoover thermostat while ensuring that the CNT remains centered in the simulation box:
@@
Here, the thermo custom command is used to print the cumulative reaction counts which are calculated by fix rxn and thus can be extracted from it. Run the simulation using LAMMPS. As the simulation progresses, polymer chains are observed forming (Fig. 49). During this reaction process, the temperature of the system remains well-controlled (Fig. 50 a), while the number of reactions, $N_{r}$, increases with time (Fig. 50 b).
@@
S.G. conceived and wrote the original online tutorials and underlying Sphinx documentation for lammpstutorials.github.io. C.A. tested the tutorials extensively, providing feedback that improved their clarity, accuracy, and usability. J.G. is the principal author of fix bond/react and type labels support in LAMMPS. He revised the tutorials to incorporate type labels and wrote Tutorial 8. A.K. developed the LAMMPS–GUI software and assisted in revising the tutorials for use with it. All authors participated in the revision and finalization of the manuscript.
@@
S.G. acknowledges funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement $\text{N}^{\circ}\;101065060$. A.K. acknowledges financial support by Sandia National Laboratories under POs 2149742 and 2407526. This work used Expanse at the San Diego Supercomputer Center through allocation MAT240081 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program.
@@ A.1 Installation
Pre-compiled versions of LAMMPS–GUI are available for Linux, macOS, and Windows on the LAMMPS GitHub Release page [21]. The Linux version is provided in two formats: as compressed tar archive (.tar.gz) and as a Flatpak bundle [71]. The macOS version is distributed as a .dmg installer image, while the Windows version comes as an executable installer package.
@@
Download the archive (e.g., LAMMPS-Linux-x86_64-GUI-22Jul2025.tar.gz) and unpack it. This will create a folder named LAMMPS–GUI containing the included commands, which can be launched directly using “./lammps-gui” or “./lmp”, for example. Adding this folder to the PATH environment variable will make these commands accessible from everywhere, without the need for the “./” prefix.
@@ A.1.2 Installing the Linux Flatpak Bundle
You have to have Flatpak support installed on Linux machine to be able to use the Flatpak bundle. Download the bundle file (e.g., LAMMPS-Linux-x86_64-GUI-22Jul2025.flatpak) and then install it using the following command:
@@
LAMMPS-Linux-x86_64-GUI-22Jul2025.flatpak
@@ A.1.3 Installing the macOS Application Bundle
After downloading the macOS app bundle image file (e.g., LAMMPS-macOS-multiarch-GUI-22Jul2025.dmg), double-click on it. In the dialog that opens drag the LAMMPS–GUI app bundle into the Applications folder. To enable command-line access, follow the instructions in the README.txt file. These macOS app-bundles contain native executables for both, Intel and Apple CPUs.
@@ A.1.4 Installing the Windows package
Download the LAMMPS–GUI installer for Windows (e.g., LAMMPS-Win10-64bit-GUI-22Jul2025.exe). Windows may warn you that the file is from an unknown developer and was downloaded from the internet. This happens because neither the installer nor the LAMMPS–GUI application (or any other included applications) have been cryptographically signed. You will need to choose to keep the file, and when launching the installer, confirm that you want to run it despite the warning.
@@ A.2 Opening, Editing, and Saving Files
Files can also be opened from the “File” menu. You can select a file through a dialog and then open it. Additionally, a history of the last five opened files is maintained, with entries to open them directly. Finally, the Ctrl-O keyboard shortcut can also be used to open a file. When integrated into a desktop environment, it is also possible to open files with a “.lmp” extension or use drag-and-drop.
@@
If the file in the editor has unsaved changes, the word “*modified*” will appear in the window title. The current input buffer can be saved by selecting “Save” or “Save As…” from the “File” menu. You can also click the “Save” icon on the left side of the status bar, or use the Ctrl-S keyboard shortcut.
@@ Appendix B Running LAMMPS on the Command-Line without the GUI
See Ref. 72 for a complete description on how to run LAMMPS.